Q) Differentiate cos 4x cos 2x with respect to `X` ?

Grade:12

2 Answers

Nandana
6 years ago
hi ,
 Ans :
        -(3 sin6x+sin2x)
Exp :
     cosA.cosB = ½ [cos(A+B) + cos(A-B)]
     cos4x.cos2x = ½[cos6x+cos2x]
    ½d/dx ([cos6x+cos2x]) =½[-6sin6x-2sin2x]
                                           = -(3 sin6x+sin2x)
                                            
                                                      – keep learning ...
 
 
 
Arun
6 years ago
cosA.cosB = ½ [cos(A+B) + cos(A-B)]cos4x.cos2x = ½[cos6x+cos2x] Now differentiating,½d/dx ([cos6x+cos2x]) =½[-6sin6x-2sin2x]= -(3 sin6x+sin2x)